OpenStudy (anonymous):

solve by substitution 3x+2y=-5 x=57-8y

OpenStudy (anonymous):

3x+2y=-5 3+-8=-5 x=1 y=-4 x=57-8y Are you supposed to plug in the Y from the first problem? If so, then: x=57-8(-4) x=57+32 x=89

hero (hero):

Something about that just looks very wrong

OpenStudy (anonymous):

I need more directions besides just the problem.

hero (hero):

The only thing that is needed is to be familiar with solving systems of equations.

hero (hero):

Solving by substitution: We're given an equation: 3x+2y=-5 We are also given an expression that is equal to x: x=57-8y Therefore, we can substitute the expression that equals x into the first equation: So: 3x + 2y = -5 becomes 3(57-8y) + 2y = -5 From there, you can simplify and solve for y.

hero (hero):

If you do that you will get: 3(57 - 8y) + 2y = -5 171 - 24y + 2y = -5 171 - 22y = -5 171 + 5 = 22y 176 = 22y 176/22 = y 8 = y

hero (hero):

Since y = 8, We can find x by substituting 8 back into the original equation: 3x + 2y = -5 3x + 2(8) = -5 3x + 16 = -5 3x = -5 - 16 3x = -21 x = -21/3 x = -7 So the solution to this system is x = -7, y = 8 or simply (-7,8).